OpenStudy (anonymous):

There are 28 students in a mathematics class. If 1/4 of the students are called to the guidance office, 1/3 of the remaining students are called to the nurse , and ,finally 1/2 of those left to go to the library , how many students remain in the classroom?

OpenStudy (mathstudent55):

28 - (1/4)28 = 28 - 7 = 21 21 - (1/3)21 = 21 - 7 = 14 14 - (1/2)14 = 14 - 7 = 7

OpenStudy (anonymous):

thanks

OpenStudy (mathstudent55):

wlcm